OpenStudy (anonymous):

Solve the following system. y = x + 4 5x + y = 16

OpenStudy (anonymous):

(2, –6) (–2, 6) (6, 2) (2, 6)

OpenStudy (anonymous):

choices :)

Directrix (directrix):

y = x + 4 5x + y = 16 5x + (x + 4) = 16 --> substituting (x + 4) for y 6x + 4 = 16 6x = 12 x = 2 ------ Okay, you find y by back substituting into y = x + 4. Post answer here.

OpenStudy (anonymous):

Is it 2,6?

Directrix (directrix):

Yes. (2, 6)

OpenStudy (anonymous):

!!! yessssss :D